OpenStudy (anonymous):

1)8 + 5n=6n what does N equal?? 2)7 + 4x + 7 =6 + 5x>>>>x=8 correct or no 3)8x + 5 = 7x + 7>>x=2 correct or no 4)4r + 12 =4 + 4r +R ......R=8 correct or not 5)1 + 7a=9+3a >>a=2 yes or no 6)11 + 7x =5+8x >>x=6 yes or no 7)7 + 4r=5r + 1+4 >>>x=??? what does X equal 8)8x +1=5+2x+8 >>x=4 yes or no 9)n +3+2n >>N=o yes or no 10)1+7m+6+8M-4m+13 .>>M=5 yes or no 11) -6.3v + 7.5= -51.09 >>v=9.3 yes or no 12) -3.7x + 1.562= -55.788 >>x=15.5 yes or no 13) -3.54 - 1.5a= -20.79 >> A=11.5 yes or no 14) 0.7x-9.1= -3.15 >>x=8.5 yes or no

jimthompson5910 (jim_thompson5910):

so far, 2) through 6) are correct

jimthompson5910 (jim_thompson5910):

8) is incorrect

OpenStudy (anonymous):

what is 8 then

jimthompson5910 (jim_thompson5910):

8x +1=5+2x+8 8x +1=2x+13 8x = 2x + 13-1 8x - 2x = 13 - 1 6x = 12 x = 12/6 x = 2

OpenStudy (anonymous):

thx

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

what about the rest are they correct

jimthompson5910 (jim_thompson5910):

9) and 10) are missing equal signs

jimthompson5910 (jim_thompson5910):

11) is correct

jimthompson5910 (jim_thompson5910):

12) is correct

OpenStudy (anonymous):

owe but there correct right

jimthompson5910 (jim_thompson5910):

13) is correct

jimthompson5910 (jim_thompson5910):

14) is correct